OpenStudy (anonymous):

Can someone help me do 3 plus the square root of 7 over 2 minus the square root of 10... It's simplifying radical expressions

OpenStudy (jacobbenvenutty):

lets go step by step 1. 3+2.64 (square root of 7) 2. Which equals 5.64 ---------------------------------------------------------------------------- 3.2-3.16 (square root of 10) 4. which equals -1.16 5. Now you should have 5.64/-1.16 (just divide those)

OpenStudy (anonymous):

thats wrong

OpenStudy (jacobbenvenutty):

.....

OpenStudy (jacobbenvenutty):

lol! sorry what did i do wrong?

OpenStudy (anonymous):

well i dont think they want the answer in decimal anyways

OpenStudy (jacobbenvenutty):

what do they want it in?

OpenStudy (anonymous):

fractions

OpenStudy (jacobbenvenutty):

So then just switch the answer to fractions

OpenStudy (anonymous):

\[is the question \frac{ \sqrt{7} }{ 2} or \sqrt{7/2} ??\]

OpenStudy (jacobbenvenutty):

True i probably should have figured that out first..... what is it @naeegee?

OpenStudy (anonymous):

yea lol he is writing in words

OpenStudy (anonymous):

I got (3 plus the square root of 7) (2 plus the square root of 10 over -6